OpenStudy (anonymous):

The graph of y=1/x^2-c has a vertical asymptote at x=3, find c. Thank you!

OpenStudy (superdavesuper):

Please clarify: is the eqn y=1/(x^2) - c or y=1/(x^2-c)

OpenStudy (anonymous):

y=1/(x^2-c)

OpenStudy (superdavesuper):

OK then y has asymptote cuz' y=1/0; so put x=3 u can solve for c

OpenStudy (anonymous):

ok thank you